### v4

# MACE-OFF: Short Range Transferable Machine Learning Force Fields for Organic Molecules

Abstract
Classical empirical force fields have dominated biomolecular simulation for over 50 years. Although widely used in drug discovery, crystal structure prediction, and biomolecular dynamics, they generally lack the accuracy and transferability required for first-principles predictive modeling. In this paper, we introduce MACE-OFF, a series of short range transferable force fields for organic molecules created using state-of-the-art machine learning technology and first-principles reference data computed with a high level of quantum mechanical theory. MACE-OFF demonstrates the remarkable capabilities of short range models by accurately predicting a wide variety of gas and condensed phase properties of molecular systems. It produces accurate, easy-to-converge dihedral torsion scans of unseen molecules, as well as reliable descriptions of molecular crystals and liquids, including quantum nuclear effects. We further demonstrate the capabilities of MACE-OFF by determining free energy surfaces in explicit solvent, as well as the folding dynamics of peptides and nanosecond simulation of a fully solvated protein. These developments enable first-principles simulations of molecular systems for the broader chemistry community at high accuracy and relatively low computational cost.
preprint: AIP/123-QED**footnotetext: These authors contributed equally.
I Introduction
Machine learning (ML) force fields have recently undergone major improvements in accuracy, robustness, and computational speed [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. They are now routinely used in materials chemistry contexts where density functional theory was previously the method of choice. In these applications, available empirical force fields, such as the embedded-atom method [14], do not provide sufficient accuracy and transferability to describe many scientifically interesting and challenging phenomena. Successful applications of ML potentials include simulation of quenching of amorphous silicon [15], determination of the phase diagrams of inorganic perovskites [16] and alloys [17], and device-scale simulation of phase-change memory materials [18].
In contrast, simulating bio-organic systems entails a different set of trade-offs, with greater emphasis on simulating large systems over long timescales. This means that empirical force fields, which sacrifice accuracy for computational speed, continue to be used routinely to study molecular liquids, crystals, biological systems, and drug-like molecules [19, 20, 21, 22].
Two alternatives to empirical force fields are available. The first is semi-empirical quantum mechanics, such as the series of extended tight-binding models [23], which represents a low-cost solution for small molecules. The method is limited by its moderate accuracy compared to quantum chemistry methods, its restriction to modelling non-periodic systems (if not in principle, but in widely used implementations), and its cubic scaling with system size.
More recently, a number of transferable machine learning force fields have also been developed for organic chemistry. By “transferable”, here we mean that these potentials are capable of generalising in system size and both chemical and conformation space sufficiently well to perform stable molecular dynamics and accurate property prediction for a wide range of molecular systems beyond those on which the model was trained. The most notable are the series of ANI [24, 25, 26, 27] and AIMNet potentials [28, 29, 8]. ANI potentials pioneered the use of local symmetry function-based feedforward neural networks [30] trained on a large dataset of organic molecular geometries [31, 32] to create transferable ML force fields. The ANI-2x model became the most widely adopted ML force field and therefore serves as one of the primary points of comparison in this paper. The ANI-2x model was recently combined with a polarizable electrostatic model [33] in a hybrid ML/MM simulation setting, and also with a neural network based dispersion correction [34]. The AIMNet models apply a message passing architecture [35], where the initial embeddings are the ANI symmetry functions. They have also extended the applicability of the models to a larger set of chemical elements, as well as to charged species. These models relax the locality assumptions by incorporating electrostatic and dispersion interactions. The PhysNet model uses a message passing architecture [35] and in addition to the semi-local terms also includes long-range electrostatic and dispersion interactions [36].
Other recent bio-organic force fields include the FENNIX model, which combines a local equivariant machine learning model with a physical long-range functional form for electrostatics and dispersion [37]. The model was trained to reproduce the CCSD(T)/CBS energies of small molecules and molecular dimers. It was shown that FENNIX can be used to run stable dynamics of liquid water, solvated alanine dipeptides, and an entire protein in the gas phase. However, further benchmarking of this model is required to assess the accuracy of the intramolecular potential outside the training set and of condensed phase molecular dynamics simulations.
Similarly, the ANA2B potential employs a short-ranged ML potential, with long-ranged, classical multipolar electrostatics, polarization and dispersion interactions [38]. Although this long-ranged model shows promising accuracy for condensed phase properties and crystal structure ranking, its accuracy and computational performance has not yet been demonstrated for larger biomolecules.
The GEMS model [39], built on the SpookyNet architecture [40], is another recent ML force field for biomolecular simulations. Whilst a transferable version of this potential is capable of producing stable dynamics for challenging biomolecular systems, certain properties, for example the folding dynamics of small peptides, require the addition of system-specific training configurations in order to accurately reproduce experimental properties. More recently, the transferable SO3LR potential was introduced by the same group of researchers [41]. This is based on the equivariant SO3krates architecture and also includes analytic dispersion and electrostatic long range contributions. Whilst this potential is capable of fast and stable simulations of biomolecular systems, we are not yet aware of more extensive quantitative benchmarking on off-equilibrium energetics or thermodynamic properties.
In this paper, we introduce a family of purely local (i.e. short range), transferable bio-organic machine learning force fields which we refer to collectively as MACE-OFF. The force fields are parameterized for the ten most important chemical elements for organic chemistry: H, C, N, O, F, P, S, Cl, Br, and I. They are capable of accurately describing intra- and intermolecular interactions of neutral closed-shell systems. This enables the simulation of a wide range of chemical systems, from molecular liquids and crystals to drug-like molecules and biopolymers.
The models were validated on a number of tasks, including the prediction of small molecule torsion barriers, geometry optimizations, calculations of lattice parameters and enthalpies of formation of molecular crystals, and the calculation of the Raman spectra of molecular crystals, including nuclear quantum effects. We have also validated the models for predicting the densities and heats of vaporization of a range of molecular liquids. In particular, we study how well MACE-OFF reproduces the fundamental properties of water, including density and radial distribution functions. To further showcase the capabilities of the model, we computed the free energy surface and J-coupling parameters of alanine tripeptide (Ala3) in vacuum and explicit water, simulated the folding of Ala15 at different temperatures, and carried out an all-atom simulation of the protein crambin in explicit water (18 000 atoms). Finally, we tested the computational speed of the current implementation in the LAMMPS and OpenMM simulation packages, demonstrating both strong and weak scaling.
While truly general purpose force fields for biomolecular modeling require explicit treatment of long-range Coulomb interactions, the purpose of this paper is to push the capabilities of short-range potentials as far as they can go, in part as preparation for inclusion of explicit polarizable electrostatic interactions in the future.
II Methods
II.1 The MACE architecture
The MACE model [2] is a force field that maps the positions and chemical elements of atoms to the system’s potential energy. Linear scaling with system size is achieved by decomposing the total energy into atomic site energies. First, a graph is defined by connecting two nodes (atoms) by an edge if they are in each other’s local environment. The local environment $\mathcal{N}(i)$ is the set of all atoms $j$ around the central atom $i$ for which $\|\bm{r}_{ij}\|\leq r_{\text{cut}}$, where $\bm{r}_{ij}$ is the vector from atom $i$ to atom $j$ and $r_{\text{cut}}$ is the cutoff hyperparameter. The array of features of node $i$ is denoted by ${\bm{h}}_{i}^{(t)}$ and is expressed in the spherical harmonic basis, with its elements indexed by $l$ and $m$. The superscript in $t$ indicates the iteration steps (corresponding to “layers” of message passing in the language of graph neural networks). All MACE-OFF models are made up of two layers. The ${\bm{h}}^{(t)}_{i}$ features depend on the chemical environment of the atoms for $t>0$.
The node features ${\bm{h}}^{(0)}_{i}$ are initialized as a (learnable) embedding of the chemical elements with atomic numbers $z_{i}$ into $k$ channels:
$h_{i,k00}^{(0)}=\sum_{z}W_{kz}\delta_{zz_{i}}$ (1)
The superscript $(0)$ in this case indicates the initial (0-th) layer. This type of mapping has been widely applied to graph neural networks [42, 43, 44] and other models [45, 46].
Next, for each atom, the features of its neighbours are combined with the interatomic displacement vectors, $\bm{r}_{ij}$, to form the one-particle basis $\phi_{ij,k\eta_{1}l_{3}m_{3}}^{(t)}$. The radial distance is used as an input into a learnable radial function $R(r_{ij})$ with several outputs that correspond to the different ways in which the displacement vector and the node features can be combined while preserving equivariance [47]:
$\phi_{ij,k\eta_{1}l_{3}m_{3}}^{(t)}$ $=\sum_{l_{1}l_{2}m_{1}m_{2}}C_{\eta_{1},l_{1}m_{1}l_{2}m_{2}}^{l_{3}m_{3}}R_{k\eta_{1}l_{1}l_{2}l_{3}}^{(t)}(r_{ij})\,\,\times$
×$\qquad\qquad\times Y^{m_{1}}_{l_{1}}(\bm{\hat{r}}_{ij})h^{(t)}_{j,kl_{2}m_{2}}$ (2)
where $Y_{l}^{m}$ are the spherical harmonics, and $C_{\eta_{1},l_{1}m_{1}l_{2}m_{2}}^{l_{3}m_{3}}$ denotes the Clebsch-Gordan coefficients. There are multiple ways of constructing an equivariant combination with a given symmetry $(l_{3},m_{3})$, and these multiplicities are enumerated by the index $\eta_{1}$ [48].
The one-particle basis $\phi$ is summed over the neighborhood and the $k$ channels are mixed together with learnable weights to form the permutation invariant atomic basis $A_{i}$:
$A_{i,kl_{3}m_{3}}^{(t)}=\sum_{\tilde{k},\eta_{1}}W_{k\tilde{k}\eta_{1}l_{3}}^{(t)}\sum_{j\in\mathcal{N}(i)}\phi_{ij,\tilde{k}\eta_{1}l_{3}m_{3}}^{(t)}$ (3)
Higher-order (many-body) symmetric features are created on each atom $i$ by taking tensor products of the atomic basis, $A_{i}$, with itself $\nu$ times, resulting in the “product basis”. The product basis is then contracted with the generalized Clebsch-Gordan coefficients, $\mathcal{C}^{LM}_{\eta_{\nu}\bm{l}\bm{m}}$, to obtain the equivariant many-body basis, $\bm{B}_{i}$ [48],
${\bm{B}}^{(t),\nu}_{i,\eta_{\nu}kLM}=\sum_{{\bm{l}}{\bm{m}}}\mathcal{C}^{LM}_{\eta_{\nu}\bm{l}\bm{m}}\prod_{\xi=1}^{\nu}A_{i,kl_{\xi}m_{\xi}}^{(t)}$ (4)
where bold $\bm{l}\bm{m}$ denotes the $\nu$-tuple of $l$ and $m$ values and similarly to Eq. (II.1), $\eta_{\nu}$ enumerates the number of possible couplings to create the features with equivariance $LM$. The maximum body-order is controlled by the parameter $\nu$ and is fixed at 3 (corresponding to 4-body terms, which includes the central atom) for all MACE-OFF models in this work.
Finally, a “message” $m_{i}$ is created on each atom as a learnable linear combination of the equivariant many-body features:
$m_{i,kLM}^{(t)}=\sum_{\nu}\sum_{\eta_{\nu}}W_{z_{i}\eta_{\nu}kL}^{(t),\nu}{\bm{B}}^{(t),\nu}_{i,\eta_{\nu}kLM}$ (5)
The recursive update of the node features ($t:0\rightarrow 2$) is obtained by adding the message to the atoms’ features from the previous iteration, with weights that depend on the chemical element ($z_{i}$):
$h^{(t+1)}_{i,kLM}=\sum_{\tilde{k}}W_{kL,\tilde{k}}^{(t)}m_{i,\tilde{k}LM}^{(t)}+\sum_{\tilde{k}}W_{kz_{i}L,\tilde{k}}^{(t)}h^{(t)}_{i,\tilde{k}LM}$ (6)
Since the initial node features ${\bm{h}}^{(0)}$ are only dependent on the chemical element of the atom, the second term of Eq. (6) is not included in the first layer. This makes it possible to set the energy of the isolated atoms (that is, those with no neighbours) exactly [49], which is often desirable [50, 37]. The MACE models in this work have an effective receptive field of $2\times r_{\text{cut}}$ due to the two layers of message passing.
The site energy is a sum of read-out functions applied to node features from the first and second layers. The read-out function is defined as a linear combination of rotationally invariant node features for the first layer, and as multi-layer perceptron (MLP) for the second layer.
$E_{i}=\sum_{t=1}^{2}E_{i}^{(t)}=\sum_{t=1}^{2}\mathcal{R}^{(t)}\left({\bm{h}}_{i}^{(t)}\right),$ (7)
$\mathcal{R}^{(t)}\left(\bm{h}_{i}^{(t)}\right)=\begin{cases}\sum_{k}W^{(t)}_{k}h^{(t)}_{i,k00}&\text{for}\;\;t=1\\[13.0pt] {\rm MLP}\left(\left\{h^{(t)}_{i,k00}\right\}_{k}\right)&\text{for}\;\;t=2\end{cases}$ (8)
The forces and stresses on the atoms are calculated by taking analytical derivatives of the total potential energy with respect to the positions of the atoms and the components of the deformation tensor in the periodic setting, as is usual.
II.2 Training data
[…]
Table 1: Summary of training and test sets. The columns ‘PubChem’ to ‘Solvated Amino Acids’ correspond to the original SPICE dataset [51].
PubChem DES370K Monomers DES370K Dimers [59] Dipeptides Solvated Amino Acids Water QMugs [60] Tripeptides
Chemical elements H, C, N, O, F, P, S, Cl, Br, I H, C, N, O, F, P, S, Cl, Br, I H, C, N, O, F, P, S, Cl, Br, I H, C, N, O, S H, C, N, O, S H, O H, C, N, O, F, P, S, Cl, Br, I H, C, N, O
System size 3-50 3-22 4-34 26-60 79-96 3-150 51-90 30-69
# Train 646821 16861 263065 19773 948 1597 2748 0
# Test 33884 889 13896 1025 52 84 144 898
The SPICE dataset only contains small molecules of up to 50 atoms. To facilitate the learning of intramolecular non-bonded interactions, we augmented SPICE with larger 50–90 atom molecules randomly selected from the QMugs dataset [60]. The geometries were generated by running molecular dynamics simulations using GFN2-xTB [23] similarly to the protocol described in Ref. [61]. The energies and forces were re-evaluated at the level of QM theory used in SPICE. Finally, to obtain a better description of water, the dataset was further augmented with a number of water clusters carved out of molecular dynamics simulations of liquid water [62], with sizes of up to 50 water molecules. In addition to the 5% of each subset, part of the COMP6 tripeptide dataset [25] was also recomputed at the SPICE level of theory and used as part of the test evaluation, but not for training.
In order to understand the role additional data plays in the performance of the model, we also investigated the difference in performance due to inclusion of additional configurations released as part of the second version of SPICE [63]. The updated dataset, which we incorporate into the MACE-OFF24(M) model (see Section II.3), contains among others, an additional 208,000 configurations comprising solvated PubChem molecules and amino acid-ligand pairs. The former were sampled from molecular dynamics using the same protocol as in the original dataset, whilst the amino acid-ligand pairs were extracted from structures in the PDB. This release also included 1000 water clusters, containing up to 90 atoms. These were also added and used to complement the set of water clusters already included in the extended dataset.
II.3 Model details
The MACE model has parameters that enable systematic control of model expressivity (and subsequent accuracy) against computational cost. In this paper, we present three variants of the MACE-OFF23 model, a small, a medium, and a large one, denoted in the text as MACE-OFF23(S), MACE-OFF23(M) and MACE-OFF23(L) respectively. The models get more accurate with size, but also have an increasing computational cost. The small model is well suited for large-scale simulations, the medium model offers a good balance of speed and accuracy, and the large model is best used for small systems or when the highest possible accuracy is desirable. MACE-OFF24(M) is an update to the medium model that uses an extended 6 Å cutoff and includes additional configurations from version 2 of the SPICE dataset [63]. The hyperparameters of the models are displayed in Table 2. All models were trained using the PyTorch [64] implementation of MACE, available at https://github.com/ACEsuit/mace. More information about the training is provided in Section S1.
Table 2: Hyperparameters of the MACE-OFF models
23(S) 23(M) 23(L) 24(M)
Cutoff radius (Å) 4.5 5.0 5.0 6.0
Chemical channels 96 128 192 128
$k$ (see Eq. (1))
max L (see Eq. (5)) 0 1 2 1
SPICE version 1 1 1 2
III Results
III.1 Extended SPICE Test Set
First, we look at the pointwise errors of the energy and force predictions on a held-out test set for each of the three baseline MACE-OFF23 models. Figure 1 shows the per-atom energy, force and intermolecular force root-mean-square errors (full statistics are provided in Section S2). As the size of the model increases, the models gradually become more accurate, with the large model generally achieving errors of around 0.5–1.0 meV/atom and 15–20 meV/Å, well below the 1 kcal/mol (43 meV) chemical accuracy limit for the drug-like organic molecules studied here.
Figure 1: Test set root mean square errors (RMSE). Errors in the MACE-OFF23 models compared to the underlying DFT reference data, highlighting the relative accuracy of the three models. Bottom panels show specifically inter-molecular force errors compared to overall DFT inter-molecular force magnitudes (RMS). Note that for subsets comprising only single molecule configurations (DES370K Monomers, Dipeptides, QMugs, Tripeptides) inter-molecular contributions are expected to be zero. The slight deviation from zero arises because DFT forces do not obey translational and rotational symmetries with sufficient accuracy, while MACE models do.
The last column represents an extrapolation task, the training set contains only dipeptides, and this test set looks at tripeptides, indicating that the models are able to extrapolate to larger fragments with more complex interactions, with no loss of accuracy.
The bottom panel of Figure 1 shows specifically the intermolecular force errors, which were obtained by separating the force contributions to molecular translations and rotations (see Section S2 and Ref. [65] for details). These interactions are crucial as they underpin the thermodynamics and transport properties in the organic condensed phase. MACE-OFF23(M) and MACE-OFF23(L) both yield similar errors of around 5–15 meV/Å, which are about 1.5–3 times smaller than total force errors, and about 5–10% of the typical DFT intermolecular force magnitude (RMS). For context, relative intramolecular force errors, which are roughly equivalent to total force errors, are around 1–2%. The latter are routinely used across the literature to benchmark ML models, however, we advocate that intermolecular RMSE errors are a more appropriate metric to assess model accuracy for organic condensed phase applications. The results here demonstrate the remarkable accuracy of MACE models in predicting these subtle but crucial interactions. We will investigate how these energy and force errors translate into condensed phase physical property predictions in later Sections.
III.2 Dihedral scans
Next, we evaluate the performance of the MACE-OFF23 models on dihedral scans of drug-like molecules. This task is routinely carried out using quantum mechanical methods to create reference data for parameterizing classical empirical force field dihedral terms [66]. The task is particularly challenging, as constrained geometry optimizations can be difficult to converge if the potential energy surface is not smooth, as has been observed for example for the ANI-1x potential [67].
Figure 2: Dihedral benchmark scans. The top panel shows torsion drive data for the TorsionNet-500 dataset, which has a wide chemical diversity (five example molecules are shown). The bottom panel focuses on the torsion angle between two aromatic rings in the biaryl torsion benchmark [68] which contains 78 molecules (five examples are shown).
III.2.1 TorsionNet-500
The top panel of Figure 2 summarizes the results for the TorsionNet-500 dataset [69]. This dataset contains torsion drives of 500 different molecules, selected to cover a wide range of pharmaceutically relevant chemical space. The original dataset was reported at the B3LYP/6-31G(d) level of DFT theory. For consistency with the SPICE dataset, we recreated the torsion profiles using the DFT setting of SPICE, which uses a higher level of theory and basis set (Section S3).
The first panel shows an example of a torsion drive, indicating the complex energy profile that the MACE models are able to capture closely, including geometries far from equilibrium. The center panel shows the mean barrier height error of a number of representative models, comparing the Sage classical empirical force field [21], a semi-empirical quantum mechanical method GFN2-xTB [23], a recent transferable machine learning force field AIMNet2 [8], and the three MACE-OFF23 models. Again, systematic improvements in accuracy with MACE model size are observed, with medium and large models, in particular, achieving errors of around 0.25 kcal/mol compared to the reference method. The AIMNet2 model achieves comparable accuracy to the small MACE-OFF23 model. A similar conclusion can be drawn from the comparison of the molecular geometries by looking at the root mean squared deviation of the atomic positions averaged over the full torsion scans, as indicated by the top right panel of Figure 2. It shows that MACE-optimized geometries have a deviation of about 0.025 Å, meaning they are almost indistinguishable from DFT-optimized structures. It is important to note that the other models were trained to different levels of DFT, which might also contribute to the observed differences [70].
III.2.2 Biaryl fragments
The biaryl torsion benchmark investigates the torsional profiles of about 100 different biaryl fragments computed at the coupled cluster level of theory [68]. In these molecules, the rotatable bond is between two aromatic rings. Such chemistry frequently occurs in drug-like molecules, and the profiles are typically challenging to model accurately using empirical classical force fields. Following previous studies, we used a subset of 78 molecules to facilitate comparisons with the ANI-1ccx model, which is only parameterized for H, C, N, and O chemical elements [68].
In the bottom panel of Figure 2, we compare the results of torsion drives of empirical force fields, semi-empirical QM methods, and the ANI-1ccx machine learning force field [26] with the MACE-OFF23 models. We also compare the DFT potential energy surfaces (using the SPICE level of theory) with the published gold standard coupled cluster data.
The DFT torsion drives achieve a mean barrier height error of 0.2 kcal/mol with respect to coupled cluster data, and is therefore the best result theoretically possible for MACE-OFF23 models which were trained on the same DFT level of theory. We find that the large model comes close to this, with a mean absolute error of 0.3 kcal/mol. The medium and small MACE models have barrier height errors of 0.4 and 0.5 kcal/mol, respectively. Remarkably, the small MACE model is significantly more accurate than the next-best non-DFT methods, ANI-1ccx and GFN2-xTB, as illustrated in the bottom-center plot of Figure 2. In particular, unlike MACE-OFF23, coupled cluster reference calculations were used to parameterize the ANI-1ccx and GFN2 models. In the bottom right, we also show the cumulative error distributions to verify that the MACE-OFF23 barrier height errors are not only accurate on average but are also robust, with essentially no outliers.
III.2.3 Infrared spectroscopy of a drug molecule
To assess the accuracy of the MACE-OFF23(M) model for finite temperature properties of molecules, we study the infrared spectrum of an isolated paracetamol molecule at room temperature. Molecular dipole moments are predicted using the MACE architecture, trained on the SPICE dataset, and quantum nucelear effects are incorporated using path-integral coarse-grained simulations (PIGS) [71]. Full details are provided in Section S8. This is a stringent test, as reproducing the line positions and their relative intensities requires an accurate treatment of both intramolecular interactions and the dipole moment in out-of-equilibrium configurations. As shown in Figure 3, we estimate the classical and approximate quantum IR spectra [71], and compare them directly with experimental data at 293 K [72]. For low and intermediate frequencies (up to around 2000 cm-1), the classical and quantum spectra agree with each other and with the experimental data, allowing identification of all collective modes. For the high-frequency O–H and N–H modes, we observe that the classical spectra are blue-shifted due to the absence of anharmonic zero-point fluctuations, which typically soften these bonds. A quantum treatment of nuclear dynamics results in good agreement with the experimental data, with the exception of a net blue shift and a slight discrepancy in the relative intensities of the high-frequency stretching with respect to the lower-frequency modes, likely due to the limitations of the reference density functional.
Figure 3: Infrared (IR) spectrum of a paracetamol molecule. The classical (blue dashed) and quantum (blue solid) IR spectrum at 293 K using the MACE-OFF23(M) model is computed by propagating the system and estimating the time correlation function of the time derivative of the total dipole moment. The experimental data are taken from Ref. 72.
III.3 Molecular crystals and organic liquids
In the following, we demonstrate the ability of the MACE-OFF force fields to simulate the vibrational and thermal properties of molecular crystals. Based on relative accuracy and computational expense, in what follows we present data only for the medium MACE-OFF models. Additional experiments, where either the higher throughput of the small model or the additional accuracy of the large model are desired, are reported in the Supporting Information.
III.3.1 Lattice enthalphies
We assess the ability of the MACE-OFF23(M) model to describe the stability of molecular crystals. We computed the enthalpies of sublimation of a range of 23 representative small molecular crystals [73] following the protocol of Ref. [74] (Section S4).
[…]
Additional analyses of the relaxed unit cell vectors using the MACE-OFF23(L) model, and the powder Raman spectrum of crystalline paracetamol are presented in Sections S5 and S8.
Figure 4: Sublimation enthalphies of molecular crystals. Comparison between predicted sublimation enthalphies of the MACE-OFF23(M) and ANI models and experiment.
III.3.2 Condensed phase properties of organic liquids
Figure 5: Molecular liquids. Comparison between MACE-OFF23(M) and ANI-2x with experiment for densities (top) and heats of vaporization (bottom) of condensed phase organic liquids.
Small molecule empirical force fields are fitted to reproduce experimental condensed phase properties including densities and heats of vaporization. As such, errors on small molecular systems are expected to be well within chemical accuracy [77, 21]. For machine learning force fields, this task represents a greater challenge, since the potential is typically trained only using QM reference data for isolated molecules and molecular dimers. From these ab initio data, the models must learn the long-range interactions necessary to reproduce bulk properties without ever fitting to them.
To investigate the performance of the MACE-OFF23 models in the condensed phase, we selected a benchmark set of 109 molecules from Ref. [78], containing a representative set of functional groups relevant to medicinal chemistry and biology. For the liquid boxes containing approximately 600 atoms, MACE-OFF23(S) achieved a throughput of $2.1\times 10^{6}$ steps/day via OpenMM on a single 80GB NVIDIA A100 GPU, whilst the medium model achieved a throughput of $1.1\times 10^{6}$ steps/day. Meanwhile MACE-OFF23(L) achieved a throughput of $2.8\times 10^{5}$ steps/day on the same system, highlighting the significant overhead compared to the other two models. See section III.5 for further discussion of the computational performance.
First, we investigate the predictions of molecular liquid densities under ambient conditions using the medium MACE model shown on the top panel of Figure 5. MACE-OFF23(M) achieves a MAE of 0.09 g/cm3, indicating that it can generalize from the intermolecular interactions between molecular dimers in the training set to the condensed phase, retaining a good correlation with experiment. In contrast, the ANI-2x potential has significantly higher errors, with MAE and RMSE errors twice as large as MACE-OFF23(M) (Section S6).
The vast majority of the outliers in the density predictions can be assigned as either low boiling-point ether-containing compounds, polychlorinated hydrocarbons or dibromo-containing compounds. In the first case, this may be due to a marginal under-prediction of the boiling point, since these compounds were simulated at 10 K below their tabulated boiling points. In the latter two cases, this functional group dependence suggests these nonbonded interactions are insufficiently represented in the present training set, which contains, for example, only 84 examples of dibromo-containing dimers. This highlights the need for additional coverage of specific functional groups within SPICE (for example, the recent AIMNet2 dataset includes 20 million configurations to cover 14 chemical elements [8]).
We further investigated the performance of MACE-OFF23(M) on heats of vaporization. The predictions correlate strongly with the experimental data; however, there is a systematic offset of approximately 2 kcal/mol (Figure 5). We hypothesized that this may result from an under-prediction of the total intermolecular interactions originating from two sources: those that appear in molecular dynamics simulations outside the effective cutoff of the MACE model (compare classical simulations, which typically employ long-range corrections to the total energy [79]), and those that the model has not learned from the limited set of intermolecular interactions in the training set.
The effect of the choice of MACE-OFF model on condensed phase properties is further investigated in Section S6. In general, we observe that the small model systematically over-predicts the experimental density (MAE 0.23 g/cm3), which is consistent with its larger intermolecular force errors (Table S1). We did not observe any significant improvements in accuracy, over the medium model, using either the large model or a model trained with an extended cutoff of 6 Å. This further reinforces the hypothesis that the poorly predicted densities are a result of under-representation of the functional groups in the training set. Nevertheless, these data indicate that, given a reasonable receptive field and a sufficiently expressive model, it is possible to fit a local force field that captures the interactions required to recover experimental properties. Further improvements in accuracy are possible without resorting to fitting to experiment by increasing further the size of the training dataset and the explicit addition of long-range Coulomb interactions into the model, which we leave for future work.
III.4 Biomolecular simulations
In the following section, we benchmark the performance of the MACE-OFF models on several well-studied large-scale biomolecular systems. This particular application area is of key interest, and state-of-the-art classical protein force fields have been carefully parameterized over many decades to reproduce key quantum mechanical and thermodynamic properties. However, this extensive parameterization process makes it difficult to extend classical force fields, and overall accuracy and transferability remain fundamentally limited by their functional form. This typically becomes apparent only when each generation of advancing hardware enables sufficiently long simulations to identify issues.
III.4.1 Water structure and dynamics
A key requirement for bio-organic force fields is an accurate description of water. Several empirical water models have been developed for this purpose, the most common for all-atom biomolecular simulations being TIP3P. Figure 6 shows that MACE-OFF23(M) correctly predicts the radial distribution function (RDF) of water, performing comparably with both TIP3P and the highly accurate MB-pol water model, whose RDF is indistinguishable from experiment [80]. These interactions are modelled in a purely local way by MACE, having been trained on small water clusters containing up to 50 water molecules. This result further emphasizes the ability of the model to generalize to bulk simulations, even though the training data only contained non-periodic clusters. In contrast, ANI-2X shows a significant over-structuring of the RDF, however this is likely driven primarily by the lack of dispersion correction in the underlying quantum mechanical method used to label the training set. Additional tests of the vibrational density of states of liquid water are presented in Section S8 and show good agreement with experiment.
Figure 6: Water structure. Radial distribution functions of MACE, ANI and TIP3P models. MB-pol traces are reproduced from Ref. [80].
We further calculated the temperature dependence of the water density using NPT molecular dynamics simulations, as shown in Figure 7. This is a highly sensitive test of the accuracy of intermolecular interactions [65]. The MACE-OFF23(M) model over-estimates the density of liquid water by approximately 20%. Since there is extensive coverage of water clusters in the training set, we hypothesized that this results from a missing long-range contribution beyond the receptive field. To investigate this hypothesis, we fitted an additional model, which has an increased 6 Å layerwise cutoff compared to 5 Å. The larger receptive field leads to a predicted water density within 2% of the experimental value at room temperature, and within 5% across the full temperature range (Figure 7). We note that similar errors are reported using many dispersion-corrected DFT exchange-correlation functionals [81]. At this point in the model development, the second version of SPICE had been released, which included Solvated Pubchem and Amino Acid Ligand Pairs datasets [63]. Given the direct applicability of these training data and the importance of modelling water to biomolecular simulations, the remainder of this section uses the MACE-OFF24(M) model, which includes the additional training data and extended 6 Å cutoff (Table 2). To confirm no loss of performance on incorporation of the additional training data, we confirmed that the test set errors are in agreement with MACE-OFF23(M) (Section S2.1).
Figure 7: Water density as a function of temperature for MACE-OFF23(M), and the extended cutoff MACE-OFF24(M) model.
III.4.2 Ala3 Free Energy Surface
We first investigated the ability of MACE-OFF24(M) to reproduce the free energy surface of the Ala3 tripeptide. We observed computational performance of $9.6\times 10^{6}$ steps/day for the vacuum system, and $2.2\times 10^{6}$ steps/day for the solvated system on an NVIDIA 80GB A100 GPU via the OpenMM interface. Although this remains significantly more expensive than the MM simulation (around $150\times 10^{6}$ steps/day), it nevertheless enables sufficient sampling of the free energy surface.
Figure 8 shows the free energy surface of Ala3 in water, modelled by AMBER14SB/TIP3P and MACE-OFF24(M), respectively. Well-tempered metadynamics was used to enhance sampling along the central $\phi$ and $\psi$ backbone angles (Section S7.1). The AMBER14SB simulation identifies four local minima, corresponding to the anti-parallel $\beta$-sheet ($\phi<-120^{\circ},\psi>120^{\circ}$), a right-handed $\alpha$-helix ($\phi=-60^{\circ},\psi>-60^{\circ}$), the corresponding left-handed $\alpha$-helix ($\phi=60^{\circ},\psi>60^{\circ}$), and a polyproline II (PPII)-type structure ($\phi=-60^{\circ},\psi>120^{\circ}$). Comparing the dihedral angle distribution with NMR J-coupling constants, the AMBER14SB simulation is generally in good agreement, with a slight overpopulation of both $\alpha$-helical structures and underpopulation of the PPII mesostate [82]. We therefore use this to enable visual comparison with a well-tuned classical forcefield.
[…]
We also compare the unbiased conformational ensemble from 20 ns of NPT dynamics sampled with MACE-OFF24(M) directly to experiment by calculating 3J coupling constants from the Ramachandran distribution. Computed values are shown in Table 3 in addition to literature values for AMBER14SB, ANI-2x and experiment [82, 83, 84]. MACE-OFF24(M) accurately predicts coupling constants, and is comparable with AMBER14SB. In agreement with our previous findings, both methods significantly outperform ANI-2x, highlighting the difficulties in computing accurate dynamics in the solution phase with MLPs.
[…]
MACE-OFF24(M) 5.70(0.43) 1.64(0.18) 1.37(0.30)
ANI-2x 7.5 1.2 1.8
AMBER14SB 6.07 1.87 1.13
EXP 5.68(0.11) 2.39(0.09) 1.13(0.08)
Table 3: Summary of 3J couplings computed from dihedral distributions. AMBER14SB and experimental values are taken from Ref. [82], and ANI-2x data from Ref. [83]. Standard errors in the mean from block averaging are shown in parentheses for MACE-OFF, and experimental uncertainties are reproduced from Ref. [82]
Figure 8: Free energy surfaces of Ala3 in explicit solvent. (left) AMBER14SB/TIP3P and (right) MACE-OFF24(M).
III.4.3 Folding dynamics of Ala15
Having confirmed that MACE-OFF24(M) is capable of recovering a good approximation to the free energy surface of a small peptide, we next investigated the folding of the longer helical peptide Ala15. Because this is a significantly larger system size than the dipeptide configurations on which the model was trained, this represents a nontrivial test of the extrapolation capability of the potential, including its ability to capture complex hydrogen bonding interactions required to stabilize the secondary structure. Such simulations are believed to be particularly difficult with purely local models such as those in the MACE-OFF family.
[…]
We also measured the secondary structure of the peptide as the temperature was increased. Starting from the folded structure at 300 K, the temperature was increased in steps from 300 K to 900 K over the course of 3 ns (Figure 10). The folded structure remains stable up to around 500 K, after which point the helix unfolds and the structure is primarily a random coil.
[…]
Figure 10: Ala15 Temperature dependence. Secondary structure assignment with increasing temperature.
III.4.4 Protein simulation in explicit solvent
Finally, we investigated the simulation of a fully solvated protein with MACE-OFF. As a test case, we chose Crambin, a 42-residue threonin storage protein that contains four charged residues. This test goes significantly beyond what can typically be expected from a local force field trained on neutral species only.
Using the MACE-OFF24(M) model, we first confirmed that the force field was indeed capable of simulating a large biomolecular system containing charged residues. In particular, it has recently been shown that the FENNIX potential fails to simulate a solvated 28 residue peptide due to unphysical proton transfers involving charged residues [37]. We confirmed that the root mean square fluctuations of the protein backbone relative to the minimized structure remained below 1 Å for the duration of a 1 ns simulation, no bond breaking occurred and the secondary structure motifs remained intact (Figure S2). We obtained a throughput of $3\times 10^{5}$ steps per day on a single NVIDIA A100 80GB GPU.
We then investigated the ability of MACE-OFF24(M) to capture the key vibrational modes of the system. To this end, we prepared a simulation box containing crambin in the high-hydration state, that is, fully solvated in explicit water, for a total system size of approximately 18,000 atoms. The power spectrum was calculated from the final 125 ps of a 1.6 ns simulation recorded at 2 fs resolution, to match the simulation length and sampling frequency employed in Ref [41]. This system has also been investigated with the AMBER classical force field and several machine learned potentials, including GEMS and, more recently, SO3LR [40, 39, 9]. All three MLPs produce spectra in qualitative agreement with each other. The water peaks at 1640 cm-1 and 3200-3600 cm-1 are reproduced reasonably accurately, especially considering that no nuclear quantum effects are taken into account (Figure 11). Furthermore, in the low frequency region, two peaks are identified in the power spectrum by GEMS, SO3LR and MACE-OFF24(M). The most obvious of these occurs at 60 cm-1 and corresponds to localized internal side chain fluctuations, while a peak at 220 cm-1 is due to specific configurations of water at the protein surface.
We note that, whilst a 125 ps trajectory is sufficient to qualitatively identify the spectral features, a much longer simulation of at least several nanoseconds would be required to fully converge the THz region of the spectrum and enable full quantitative comparison with experiment, as demonstrated for the GEMS model in Ref. [39]. We note that the power spectrum computed on the final 1.3 ns simulation (discarding the initial 300 ps to equilibration) does not yield a significantly different spectrum to that of the 125 ps simulation used in Figure 11, indicating that significantly more sampling would be required to resolve the THz lineshapes sufficiently to enable experimental comparison [87].
It is also noteworthy that whilst the three MLPs compared in this section give a qualitatively similar result, their modelling choices are significantly different. Most notably, GEMS and SO3LR explicitly include long-range interactions, whilst MACE is local. Additionally, of these models, both MACE and SO3LR are universal potentials, whilst GEMS is fitted in a system-specific fashion. We mention that, in addition to the GEMS model, the authors benchmarked a variant which was trained only on small fragments, denoted GEMS*, instead of the large system-specific configurations used to fit GEMS [39]. The power spectrum produced by this model is also in qualitative agreement with the other ML potentials, however, the RMSD is reduced, indicating the protein structure is confined to small fluctuations around local minima. This model also fails to reproduce experimental observables of other biomolecular systems, for example Ala15 is not predicted to fold under the GEMS* potential, suggesting the top-down fragments are required to accurately reproduce nonbonded interactions within the GEMS family of models.
The results of sections III.3 and III.4 highlight the most significant differences between previous transferable organic ML potentials and the MACE-OFF models. Whilst there is still significant work to be done to comprehensively show improvement over classical forcefields for large biomolecular systems, these results demonstrate the capabilities of models fit ‘bottom up’ from quantum mechanical data to learn subtle intermolecular interactions and correctly predict macroscopic properties.
Figure 11: Protein in explicit solvent. Power spectrum of the high hydration state of crambin computed from the final 125 ps of a 1.6 ns simulation with MACE-OFF24(M). GEMS, AMBER and SO3LR spectra are reproduced from Ref. [41]. Black lines denote experimental peaks at 1640 cm-1 and the range 3200-3600 cm-1.
III.5 Computational performance
To evaluate the computational performance of the MACE-OFF models, we performed NVT molecular dynamics for water using both OpenMM and LAMMPS, with simulation boxes ranging from $10^{2}$ to $10^{5}$ atoms. The OpenMM results shown in Figure 12 were collected with a single NVIDIA A100 80GB GPU. The LAMMPS results shown in Section S9 highlight both CPU and GPU evaluation.
In addition to the native PyTorch implementations of the MACE-OFF23 models, we also report results for optimized implementations leveraging custom C++, CUDA, and/or Kokkos code. One optimized variant incorporates custom CUDA kernels which are packaged as a standalone, inference-only library, cuda_mace [88]. Another pure-C++ implementation, as well as a Kokkos implementation, will be made available as part of the symmetrix library [89].
In these optimized variants, the learnable radial function $R^{(t)}_{k\eta_{1}l_{1}l_{2}l_{3}}(r_{ij})$ is evaluated with cubic splines, which efficiently map $r_{ij}$ to each of the outputs of the radial MLP. Furthermore, eqs. II.1 and 3 are implemented in unified kernels that simultaneously compute the one-particle basis and perform the summation over neighbors to generate the invariant atomic basis $A_{i}$. The computation of spherical harmonics, as detailed in equation II.1 is accelerated using the sphericart [90] library, which offers a highly optimized implementation for real spherical harmonics. All linear and equivariant linear operations have been substituted with custom kernels, and the cuda_mace implementation further leverages tensor core acceleration while incorporating error-correction techniques to ensure numerical accuracy. Additionally, the computation of the product basis $\mathbf{B}_{i}$ in eq 4 is optimized using the sparsity of the generalized Clebsch-Gordon coefficients $\mathcal{C}^{LM}_{\eta_{\nu}\bm{l}\bm{m}}$, significantly accelerating the product calculations over both the body order $v$ and the coupling terms.
Figure 12: Performance in OpenMM. Performance of MACE-OFF models in OpenMM for liquid water in the NVT ensemble at 1 g/cm3 at 300 K with a 1 fs timestep on a single NVIDIA 80GB A100 GPU. The cuda_mace implementation is labeled ‘OPT’. Forces were evaluated in double precision, and integrated in single precision.
IV Outlook
In this paper, we presented the series of MACE-OFF force field models, demonstrating the broad applicability, transferability and capability of purely short-range machine learning potentials for organic (bio)molecular simulations. We have shown that the models, based on the MACE higher order equivariant message passing architecture, can improve on the pioneering ANI models [27] that served as the only widely applicable machine learning force fields for molecules for a number of years. While a full ablation study between different ML models is not possible due to a wide range of design choices made in their construction, the MACE-OFF model architecture, training data and loss function all contribute to significant improvements in both accuracy and extrapolation, combined with high computational speed.
We trained here so-called small, medium and large versions of the MACE-OFF models, with systematically improving accuracy when benchmarked against gas phase quantum mechanical data. We further investigated the accuracy of the medium models against a range of crystalline, liquid and biomolecular condensed phase data, as we found them to provide a good compromise between accuracy and computational expense. We have made all force fields available to the community, so that users can choose the model most suited to their application.
The lack of explicit long-range interactions limits the domain of applicability of the present model to neutral, non-radical and non-reactive systems. This is something that the recently published AIMNet2 model aims to address by extending the ANI models to include charged species and long-range interactions [8]. We are currently working on a next-generation MACE-OFF model that will similarly include an explicit description of charges, enabling the description of amino acids with different protonation states, charged nucleic acids, and counter-ions. This will pave the way towards obtaining an accurate quantum mechanical transferable machine learning force field for simulating the full range of biologically relevant systems.
Acknowledgements
DPK acknowledges support from AstraZeneca and the EPSRC. DJC and JH acknowledge support from a UKRI Future Leaders Fellowship (grant MR/T019654/1). V.K. acknowledges support from the Ernest Oppenheimer Early Career Fellowship, the Sydney Harvey Junior Research Fellowship, Churchill College, University of Cambridge, and UCL’s startup funds. W.C.W. acknowledges support from the EPSRC (Grant EP/V062654/1), as well as an engagement from the Virtual Institute for Scientific Software at Georgia Tech, funded by Schmidt Sciences. In particular, Ketan Bhardwaj and Dave Brownwell of the Georgia Tech Scientific Software Engineering Center contributed to the Kokkos-based MACE evaluator. JHM acknowledges support from an AstraZeneca Non-Clinical PhD studentship, and thanks Prof. Ola Engkvist, Dr Marco Klähn and Dr Graeme Robb from AstraZeneca for their helpful discussions and supervision. The authors also thank Christoph Schran for providing the carved water cluster geometries. We also thank Dr Peter Eastman and Dr Stephen Farr for their help implementing the MACE-OpenMM interface.
We are grateful for computational support from the Swiss National Supercomputing Centre under project s1209, UCL Myriad High Performance Computing Facility, the Rocket High Performance Computing service at Newcastle University, the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and EPSRC grant EP/P022561/1, the Cambridge Service for Data Driven Discovery (CSD3), obtained through a University of Cambridge EPSRC Core Equipment Award (EP/X034712/1), and Ångström AI.
Contributions
DPK conceived the project, curated the training data, trained and evaluated the models, ran parts of the torsion drive tests, the molecular crystal tests, fitted the dipole and polarisability models for water and paracetamol bulk systems, and wrote the initial draft. JHM and DPK implemented the OpenMM interface. JHM trained and evaluated the models, curated the datasets, performed experiments on molecular liquids, peptides and proteins, benchmarked the model performance in OpenMM and wrote the initial draft. JTH implemented the MACE-torsion drive interfaces, and ran the DFT and force field torsion drives. VK fitted the MACE-OFF23-$\mu$, performed the MD and PIMD simulations of water and molecular crystals and produced their IR and Raman spectra. YP performed the error analysis for the MACE-OFF23-$\mu$ model and performed MD and PIMD simulations and calculated the IR spectrum of paracetamol molecule. IB and NJB implemented the accelerated custom kernels, and the cuda_mace library is maintained by NJB. WCW implemented the MACE-LAMMPS interface, wrote the pure C++ and Kokkos implementations, and benchmarked the model performance in LAMMPS. IBM developed the intermolecular error analysis and evaluated the models. DJC and GC supervised the project. All authors discussed the results and edited the final manuscript.
Data Availability
The data used to train the models are publicly available at: https://doi.org/10.17863/CAM.107498. The torsion drive dataset is also available at: https://zenodo.org/records/11385284. The MACE-OFF series of models is available at: https://github.com/ACEsuit/mace-off/tree/main.